Calculate the momentum of a proton whose kinetic energy is 200 MeV.


1
Expert's answer
2021-04-09T11:25:32-0400

Given

"E^2 - p^2 c^2 =m^2 c^4" where E is full energy, p is momentum and m - mass,

and

"E = T+E_0 = T+mc^2"

we obtain

"(T+mc^2)^2 - p^2c^2 = m^2 c^4"

"T^2 +2Tmc^2 + m^2c^4 -p^2c^2 =m^2c^4"

"p^2c^2 = T^2 +2Tmc^2"

"pc = \sqrt{T^2+2Tmc^2}"

Let's calculate all in units of GeV, "T=200\; MeV , mc^2 =938 \; MeV".

"pc = \sqrt{200^2 + 2\cdot 200 \cdot 938}=\sqrt{415\,200}=644.36 \; MeV"


Answer: p = 644.36 MeV/c.